# Analytical exciton energies in monolayer transition-metal dichalcogenides

### v2

@@ 2 Department of Physics, University of South Florida, Tampa, Florida 33620, USA
S-I Algebraic form of the Schrödinger equation for an exciton in monolayer transition-metal dichalcogenides
S-I.1 The Schrödinger equation for an electron-hole system via the Levi-Civita transformation
@@
where $\alpha=r_{0}/\kappa a_{0}^{*}$. We note that most studies use the Keldysh potential expressed via Struve and Bessel functions, which are suitable only for numerical calculations.
@@
S-I.2 Algebraic formalism via annihilation and creation operators
We rewrite the Schrödinger equation (S-3) in the form
@@
Equation (S-6) is similar to those describing a two-dimensional anharmonic oscillator, suggesting we use its algebraic representation via annihilation and creation operators for convenience in calculations. We then use the harmonic osscilator’s wave functions to establish a basis set and calculate all necessary matrix elements by the algebraic method. More details of this algebraic approach can be found in monograph [22].
@@
Here, $\omega$ in the above definition is usually an angular frequency of the harmonic oscillator. However, we can freely choose its value to manipulate the converge rate of the solutions, see [22] and references therein. Then, we use a canonical transform to get operators
@@
Basis set – The Hamiltonian of two-dimentional isotropic harmonic oscillator associated with the annihilation and creation operators (S-I.2) has the algebraic form $\hat{H}_{osc}=\omega({\hat{a}}^{+}{\hat{a}}+{\hat{b}}^{+}{\hat{b}}+1)$. We will choose the eigenvectors of this Hamiltonian to be a basis set for our resolving process considering the Schrödinger equation (S-4) describing a two-dimensional anharmonic oscillator in the $(u,v)$- space. In addition, because the angular momentum conservation of excitons in monolayer TMDCs, we also require the basis set functions to be eigenfunctions of the operator ${\hat{l}}_{z}$ (Eq.(S-13)). As a results, we obtain
@@
as a basis set, where the vacuum state $|0(\omega)\rangle$ sastifies the equations:
@@
Considering the algebraic form (S-13), we can verify that ${\hat{l}}_{z}\,{|k,m\rangle}=m\,{|k,m\rangle}$ with $m=0,\pm 1,\pm 2,...$. Therefore the running index $m$ of the basis set is fixed by the magnetic quantum number of the system. The other running index $k$ is integer with the values $k\geq|m|$. By using the commutation relations (S-12), (S-14) and equations (S-18), we can also certify the orthonormality of the basis set functions (S-17), i.e. $\langle k,m|k^{\prime},m^{\prime}\rangle=\delta_{k,k^{\prime}}\delta_{m,m^{\prime}},$ where $\delta_{ij}$ is the Kronecker delta.
Matrix elements – Commutation relations (S-12), (S-14) and equations (S-18) are useful for constructing algebraic calculation technique based on the annihilation and creation operators. This kind of algebraic techique is given in textbooks on quantum mechanics and monograph [22]. For all the operators in the Schrödinger equation (S-4), matrix elements with repect to the basis set (S-17) can be calculated and formulated analytically in Ref. [15]. In the present paper, we consider only $s$-states ($m=0$) so the basis set becomes
@@
that means operators $\hat{N}$, $\hat{M}$, and $\hat{M}^{+}$ creat a close algebra. Following this procedure given on page 232-233 of Ref. [22], we have the formula
@@
$h_{j}$ $=$ $\langle j|\hat{H}|j\rangle=-\frac{1}{4}T_{jj}+V^{\text{\tiny K}}_{jj}=\frac{\omega}{4}(2j+1)-\frac{2}{\omega\alpha}\left[(2j+1)u_{jj}+ju_{j,j-1}+(j+1)u_{j,j+1}\right],$
@@
$=$ $\frac{\omega}{4}\left(k\delta_{j,k-1}+j\delta_{j,k+1}\right)-\frac{2}{\omega\alpha}\left[(2k+1)u_{j,k}+ku_{j,k-1}+(k+1)u_{j,k+1}\right],$
@@
where $r_{0}^{*}=r_{0}/a_{0}^{*}$ is a dimensionless screening length (in an unit of effective Bohr radius scaled by the dielectric constant). There is a free parameter $\lambda$ in energies (S-57) which we need to define. In principle, numerical exact solutions do not depend on this parameter, but approximate solutions are dependent on the choice of this parameter. Previous studies [22, 15] showed that choosing the free parameter around the solution of equation
@@
is optimal, leading to accurate energies within some lowlying corrections. We choose the parameter $\lambda$ from Eq. (S-58) and put it to expression (S-57) to get energy in the zer0-order approximation. By numerical calculations with formulas (S-51) and (S-53), we obtain also exciton energies in second- and third-order approximations. Material parameters are picked up from Table S-I. Results in Tables S-II, S-III, S-IV, and S-V for monolayers WSe2, WS2, MoSe2, MoS2 encapsulated in hBN slabs demonstrate extremely high precision, within 1.0 meV in second-order approximation and 0.5 meV in third-order approximation.
@@
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{\text{num.}}$ [15]
$1s$215.29 -219.38 -220.00 -220.18
$2s$53.09 -49.52 -51.66 -51.81
$3s$23.01 -21.94 -22.18 -22.46
$4s$12.71 -12.28 -12.27 -12.44
$5s$8.03 -7.82 -7.82 -7.89
S-III Analytical expression for $s$-state exciton energies
@@
Polynomial $P_{1}(\delta\lambda)$ in Eq. (S-67) is approximately equal to one because of the smallness of the derivation $\delta\lambda$ ($0.062$ WSe2, $0.128$ WS2, $-0.173$ MoSe2, and $0.0003$ MoS2). To evaluate this polynomial, we show in Tab. S-VI its values for different monolayer TMDCs. We can see that, the polynomial has almost the value of 1.0 for MoS2 and can be truncated at $\delta\lambda^{0}$. For monolayers WSe2, this polynomial is equal to 0.99 and truncated at $\delta\lambda^{1}$. The same for monolayer WS2, the polynomial has the value of 0.987 but truncated at $\delta\lambda^{3}$. The worst result is for monolayer MoSe2 where the polynomial has the value of 1.066 and is truncated at $\delta\lambda^{5}$.
@@
where $P_{1}$ is a polynomial whose values is around 1.0 as provided in Tab. S-VI. From the other hand, the energy spectra of two-dimensional hydrogen atom has the form
@@
Thus, the exciton energy now become
@@
From Table S-VII, we can see that polynomial $P^{*}_{1}$ truncated at $\sim\delta\lambda^{4}$ is accurate within a precision of 0.2%. Table S-VIII gives exciton energies for the $1s$ state calculated by analytical formula (S-73) compared with the exact numerical solutions. The accuracy of the formula is very hight, less than 1.0%.
@@
$E_{ns}^{Molas}$ $=$ $-\frac{\text{Ry}^{*}}{(n+\delta)^{2}}.$ (S-115)
Our formula is similar to the one of Molas with $\delta=-0.5+0.479*r_{0}^{*}/\kappa$ and $\text{Ry}^{*}(modify)={\text{Ry}}^{*}\times P_{n}$. For monolayer WSe2 in bHN, Molas’s formula is $E_{ns}^{Molas}=134{\text{meV}}/(n-0.099)^{2}$ while our formula is rewritten as $E_{ns}=134.5{\text{meV}}/(1-0.11)^{2}$. Notably, in our formula the effect of screening the charge is presented via the correction $P_{n}^{*}$.
@@
We use experimental data of Refs. [9, 10, 8] to test analytical formula (S-115). Some data are provided numericallly in the works, other data we extract from the figures of the works. Theoretically, we calculate exciton energies for monolayer Wse2, WS2, MoSe2, and MoS2 by formula (S-115) with material parameters picked up from Refs. [10] and [15]. For MoTe2, Ref. [10] give only two energy values: $E_{b}=177\pm 3$ meV and transition energy $2s-1s$ 124 meV. Material parameters given in Ref. [10]: $\mu=0.36\pm m_{e}$, $r_{0}=6.4\pm 0.3$ nm are in a wide interval of encertainty while $\kappa=4.4$ is picked up from other works. Choosing the material parameters so that our theory best fit with the experimental data we have $\mu=0.37m_{e}$, $\kappa=4.3$, and $r_{0}=6.1$ nm. We show the experimental data and theoretical exciton energies in Figs. S-1, S-2, and S-3, which demonstrate the high consistency of our theory with the experiments. Experimental data of Liu et al. [8] are not shown in the figures because they are almost identified with the ones of Ref. Stier2018. Numerically, they match with our theory. Indeed, for monolayer WSe2 , $E_{1s}=-172$ meV, $E_{2s}=-41$ meV, $E_{3s}=-20$ meV, $E_{4s}=-11$ meV. They lead to the transition energies $2s-1s$ 131 meV, $3s-2s$ 21 meV, $4s-2s$ 30 meV. Our theory give -168.38, -38.34, -16.32, and - 8.97 meV for energies, and $2s-1s$ 130.04 meV, $3s-2s$ 22.02 meV, $4s-2s$ 29.37 meV for transition energies.

### v3

@@ 2 Department of Physics, University of South Florida, Tampa, Florida 33620, USA
S-I Analytical matrix elements
S-I.1 The Schrödinger equation for an exciton in monolayer TMDCs via the Levi-Civita transformation
@@
where $\alpha=r_{0}/\kappa a_{0}^{*}$ with the screening length $r_{0}$. We note that most studies use the Keldysh potential expressed via Struve and Bessel functions, which are suitable only for numerical calculations.
@@
S-I.2 Algebraic formalism with annihilation and creation operators
We rewrite the Schrödinger equation (S-4) in the form
@@
Equation (S-7) is similar to those describing a two-dimensional anharmonic oscillator, suggesting we use its algebraic representation via annihilation and creation operators for convenience in calculations. Using the algebraic method, we then use the harmonic oscillator’s wave functions to establish a basis set and calculate all necessary matrix elements. More details of this algebraic approach can be found in monograph [24].
@@
Here, $\omega$ in the above definition is usually an angular frequency of the harmonic oscillator. However, we can freely choose its value to manipulate the converge rate of the solutions, see [24] and references therein. Then, we use a canonical transformation to obtain operators
@@
Basis set – The Hamiltonian of two-dimentional isotropic harmonic oscillator associated with the annihilation and creation operators (S-I.2) has the algebraic form $\hat{H}_{\text{osc}}=\omega({\hat{a}}^{+}{\hat{a}}+{\hat{b}}^{+}{\hat{b}}+1)$. We will choose the eigenvectors of this Hamiltonian to be a basis set for our resolving process considering the Schrödinger equation (S-4) describing a two-dimensional anharmonic oscillator in the $(u,v)$- space. In addition, because the angular momentum conservation of excitons in monolayer TMDCs, we also require the basis set functions to be eigenfunctions of the operator ${\hat{l}}_{z}$ (Eq.(S-13)). As a results, we obtain
@@
as a basis set, where the vacuum state $|0(\omega)\rangle$ satisfies the equations:
@@
Considering the algebraic form (S-13), we can verify that ${\hat{l}}_{z}\,{|k,m\rangle}=m\,{|k,m\rangle}$ with $m=0,\pm 1,\pm 2,...$. Therefore the running index $m$ of the basis set is fixed by the magnetic quantum number of the system. The other running index $k$ is an integer with the values $k\geq|m|$. By using the commutation relations (S-12), (S-14) and equations (S-18), we can also certify the orthonormality of the basis set functions (S-17), i.e. $\langle k,m|k^{\prime},m^{\prime}\rangle=\delta_{k,k^{\prime}}\delta_{m,m^{\prime}},$ where $\delta_{ij}$ is the Kronecker delta.
Matrix elements – Commutation relations (S-12), (S-14), and equations (S-18) are useful for constructing algebraic calculation techniques based on the annihilation and creation operators. This kind of algebraic technique is given in textbooks on quantum mechanics and monograph [24]. For all the operators in the Schrödinger equation (S-4), matrix elements with respect to the basis set (S-17) can be calculated and formulated analytically in Ref. [16]. In the present paper, we consider only $s$-states ($m=0$) so the basis set becomes
@@
that means operators $\hat{N}$, $\hat{M}$, and $\hat{M}^{+}$ creat a close algebra. Following this procedure given on page 232–233 of Ref. [24], we have the formula
@@
$h_{j}$ $=$ $\langle j|\hat{H}|j\rangle=-\frac{1}{4}T_{jj}+V^{\text{\tiny K}}_{jj}=\frac{\omega}{4}(2j+1)-\frac{2\kappa}{\omega r_{0}^{*}}\left[(2j+1)u_{jj}+ju_{j,j-1}+(j+1)u_{j,j+1}\right],$
@@
$=$ $\frac{\omega}{4}\left(k\delta_{j,k-1}+j\delta_{j,k+1}\right)-\frac{2\kappa}{\omega r_{0}^{*}}\left[(2k+1)u_{j,k}+ku_{j,k-1}+(k+1)u_{j,k+1}\right],$
@@
where $r_{0}^{*}=r_{0}/a_{0}^{*}$ is a dimensionless screening length (in a unit of effective Bohr radius scaled by the dielectric constant). There is a free parameter $\lambda$ in energies (S-57), which we need to define. In principle, numerical exact solutions do not depend on this parameter, but approximate solutions are dependent on the choice of this parameter. Previous studies [24, 16] showed that choosing the free parameter around the solution of the equation
@@
is optimal, leading to accurate energies within some low-lying corrections. We choose the parameter $\lambda$ from Eq. (S-58) and put it to expression (S-57) to get energy in the zer0-order approximation. By numerical calculations with formulas (S-51) and (S-53), we also obtain exciton energies in second- and third-order approximations. Material parameters are picked up from Table S-I. Results in Tables S-II, S-III, S-IV, and S-V for monolayers WSe2, WS2, MoSe2, MoS2 encapsulated in hBN slabs demonstrate extremely high precision, within 1.0 meV in second-order approximation and 0.5 meV in third-order approximation.
@@
$E^{(0)}$ $E^{(2)}$ $E^{(3)}$ $E_{\text{num}}$ [16]
$1s$ $-215.29$ $-219.38$ $-220.00$ $-220.18$
$2s$ $-53.09$ $-49.52$ $-51.66$ $-51.81$
$3s$ $-23.01$ $-21.94$ $-22.18$ $-22.46$
$4s$ $-12.71$ $-12.28$ $-12.27$ $-12.44$
$5s$ $-8.03$ $-7.82$ $-7.82$ $-7.89$
S-III Derivation of analytical expression for $s$-state exciton energies
@@
Polynomial $P_{1}(\delta\lambda)$ in Eq. (S-67) is approximately equal to one because of the smallness of the derivation $\delta\lambda$ ($0.062$ WSe2, $0.128$ WS2, $-0.173$ MoSe2, and $0.0003$ MoS2). To evaluate this polynomial, we show in Tab. S-VI its values for different monolayer TMDCs. We can see that the polynomial has almost the value of 1.0 for MoS2 and can be truncated at $\delta\lambda^{0}$. For monolayers WSe2, this polynomial is equal to 0.99 and truncated at $\delta\lambda^{1}$. The same for monolayer WS2, the polynomial has the value of 0.987 but is truncated at $\delta\lambda^{3}$. The worst result is for monolayer MoSe2 where the polynomial has the value of 1.066 and is truncated at $\delta\lambda^{5}$.
@@
where $P_{1}$ is a polynomial whose value is around 1.0, as provided in Tab. S-VI. On the other hand, the energy spectra of the two-dimensional hydrogen atom have the form
@@
Thus, the exciton energy now becomes
@@
Table S-VII shows that polynomial $P^{*}_{1}$ truncated at $\sim\delta\lambda^{4}$ is accurate within a precision of 0.2%. Table S-VIII gives exciton energies for the $1s$ state calculated by the analytical formula (S-73) compared with the exact numerical solutions. The accuracy of the formula is very high, less than 1.0%.
@@
$E_{ns}^{\text{Molas}}$ $=$ $-\frac{\text{Ry}^{*}}{(n+\delta)^{2}}.$ (S-115)
Our formula is similar to the one of Molas with $\delta=-0.5+0.479*r_{0}^{*}/\kappa$ and $\text{Ry}^{*}(\text{modify})={\text{Ry}}^{*}\times P_{n}$. For monolayer WSe2 in bHN, Molas’s formula is $E_{ns}^{\text{Molas}}=134{\text{meV}}/(n-0.099)^{2}$ while our formula is rewritten as $E_{ns}=134.5{\text{meV}}/(1-0.11)^{2}$. Notably, in our formula, the effect of screening the charge is presented via the correction $P_{n}^{*}$.
@@
We use experimental data of Refs. [9, 10, 8] to test analytical formula (S-115). Some data are provided numerically in these works, and we extract other data from the figures of the works. Theoretically, we calculate exciton energies for monolayer Wse2, WS2, MoSe2, and MoS2 by formula (S-115) with material parameters picked up from Refs. [10] and [16]. For MoTe2, Ref. [10] give only two energy values: $E_{b}=177\pm 3$ meV and transition energy $2s-1s$ 124 meV. Material parameters given in Ref. [10]: $\mu=0.36\pm m_{e}$, $r_{0}=6.4\pm 0.3$ nm are in a wide interval of uncertainty while $\kappa=4.4$ is picked up from other works. Choosing the material parameters so that our theory best fits with the experimental data, we have $\mu=0.37m_{e}$, $\kappa=4.3$, and $r_{0}=6.1$ nm. We show the experimental data and theoretical exciton energies in Figs. S-1, S-2, and S-3, which demonstrate the high consistency of our theory with the experiments. Experimental data of Liu et al. [8] are not shown in the figures because they are almost identified with the ones of Ref. [9]. Numerically, they match our theory. Indeed, for monolayer WSe2 , $E_{1s}=-172$ meV, $E_{2s}=-41$ meV, $E_{3s}=-20$ meV, $E_{4s}=-11$ meV. They lead to the transition energies $2s-1s$ 131 meV, $3s-2s$ 21 meV, $4s-2s$ 30 meV. Our theory give $-168.38$, $-38.34$, $-16.32$, and $-8.97$ meV for energies, and $2s-1s$ 130.04 meV, $3s-2s$ 22.02 meV, $4s-2s$ 29.37 meV for transition energies.